A bank charges its customers a general fee of $6 per month plus an additional fee of 10 cents per check written from a checking account. Over the past nine months, a customer was charged $62 in service fees. How many checks did the customer write during that period?
Answer by checkley77(12569)

9*6+.10x=62
54+.10x=62
.10x=62-54
.10x=8
x=8/.10
x=80 checks were written during the 9 months.
Proof:
9*6+.10*80=62
54+8=62
62=62